# Chapter 0 Hadron Production Processes

### v1

@@ 1 Aspects of QCD
The seemingly simple picture of hadrons emerging at first sight from QCD is in fact deceiving because the complex field-theoretical dynamics lead in fact to a large variety of spectroscopic configurations. The basic constituents, quarks and gluons, are interacting in a manifold manner which will lead to a variety of intrinsic configurations forming the inner confined core of a hadron. A selection of expected elementary configurations is displayed in Fig. 8. The core region of a physical hadron will contain a mixture of those configurations, i.e., hadron wave functions are typically quantum mechanical superpositions of multi-quark, quark-gluon, gluon-gluon, etc. components - all embedded on top of a highly dynamical vacuum. The core region, albeit confined to an incredibly small volume of about $(0.5\times 10^{-45})~{}\mathrm{m^{3}}$, is only part of the complete hadron. The quark-gluon core polarizes the vacuum and, as a consequence, is surrounded by a cloud of virtual particles, mainly consisting of (pairs of) pions and other mesons.
@@
Asymptotic freedom on the one side and quark confinement on the other side are defining parts of strong interaction physics. Gluon self-interactions and effects from the spontaneously broken chiral symmetry are important ingredients of constituent quark and hadron masses. The highly non-linear nature of QCD inhibits analytical or perturbative solutions for most parts of the accessible energy region, except for the highest energies as, e.g., reachable at the LHC at CERN. The mathematical complexities of QCD at energies relevant for hadron structure, spectroscopy, and interactions of hadrons can be treated only numerically. LQCD is the method of choice, being used successfully to explore mesons and baryons by QCD principles.
@@ 1 QCD and Phenomenology
While the high energy sector of QCD is well understood on the basis of perturbative QCD, the low-energy limit is still a field with many open questions, reflecting the highly nontrivial task of handling a non-perturbative quantum field theory. Formally, the QCD path integral is the proper starting point. As discussed elsewhere in this encyclopedia, two alternative numerical schemes are currently pursued: In LQCD, a set of coupled correlation functions is derived from the path integral and propagated on a Euclidean (named after Euclid (about 325 — 265 BC), an ancient Greek mathematician) space-time grid where masses are determined by the numerically produced pole structures [57]. One of the first successful LQCD descriptions of the mass spectrum of the octet mesons and the octet and decuplet baryons, respectively, was achieved by the Budapest-Wuppertal Collaboration [58, 59]. The mass spectrum is displayed in Fig. 9. Functional methods derive sets of coupled Dyson-Schwinger equations (DSE) (named after Freeman John Dyson (1923-2020) and Julian Schwinger) by performing variational functional derivatives of the path integral [60].
@@
While LQCD and DSE approach hadron spectroscopy from the inner sector of the quark core of hadrons, effective field theories and phenomenological reaction models put their focus on the properties reflecting the nature of hadrons as dynamically generated composite states. Compositeness is another aspect of the dualism hidden in baryons and mesons as visible in the observed resonances in scattering and production experiments. Until today, phenomenological approaches based on an underlying covariant field theory are an important source of spectroscopic information on the quantum numbers of resonances, their formation, and their decay. The models are apt for large-scale numerical coupled channels calculations utilizing the partial wave formalism. Meson production on the nucleon is described through the excitation of baryon resonances in photo- and meson-induced reactions on the nucleon. Baryon resonances are identified through their traces left as complex energy poles in the partial wave scattering matrix elements, which in many cases are strongly affected by coupled channel effects.
@@
At masses above the pion-nucleon threshold, baryon resonances have to be considered as superpositions of molecular-like, loosely bound, or unbound meson-nucleon configurations. They are in competition with the genuine QCD-type confined quark core configurations, illustrated in Fig. 8. In addition, the QCD core is shielded from the exterior by a virtual $q\bar{q}$ cloud of transient mesons. Very likely, also glue-ball-like configurations will contribute as indicated by the large differences between the current and the constituent quark masses. A given resonance will be composed of an arbitrary mixture of two building blocks. In some cases, the quark-type configurations will dominate; in others, the molecular configuration may prevail. An intensively and controversially debated candidate for the latter type of states is the $\Lambda^{\ast}(1405)$ ($S=-1$) resonance, which falls out of the systematics of quark models but is characterized by a pronounced pole structure in the complex plane. The chiral unitary model [61] predicted a two-pole structure from a superposition of $\pi\Sigma$ and $\bar{K}N$ components. Years later, improved experimental and theoretical studies concluded that $\Lambda^{\ast}(1405)$ is dominantly a $\bar{K}N$ composite; however, without fully excluding $\pi\Sigma$ admixtures [62].
@@ 5 The Constituent Quark Model and Hadron Spectroscopy
The discovery of resonance phenomena in the early pion-nucleon scattering experiments cast strong doubts on the initial assumption that nucleons were structure-less elementary point particles, as was assumed - and since then confirmed - for the electron.
Table 1: The mass spectrum of constituent quarks. the masses are derived from the data od meson and baryon masses such that the experimental masses are reproduced by the sum of constituent quark masses. The quark generations are indicated by separating lines. For further details see Ref. [63].
@@
Once QCD had established quarks and gluons as the constituents of hadrons, the interest in nucleon spectroscopy was inevitable. The CQM, going back to Isgur and Karl [64, 65, 66, 67, 68, 69, 70, 71], was an important step forward to understand and predict hadron spectra on phenomenological grounds. Mesons are described by quark-antiquark ($q\bar{q}$) configurations and baryons as states of three valence quarks ($qqq$). The gluon component in hadron masses - actually accounting for the major part of the masses - is taken into account by assigning to the quarks effective constituent masses which are derived empirically from data. The constituent quarks interact by a Coulomb-like static potential [^10] Charles-Augustin de Coulomb (1736-1806) derived first the $1/r^{2}$ behavior of forces between electric charges, corresponding to a potential $\sim 1/r$., modeling gluon exchange, and are enclosed by a confinement potential, typically chosen as a power law in the radial coordinate. Spin- and flavor-dependent residual interactions are used to describe the fine structure of hadrons. Modern versions of the Cornell potential, originally introduced in the early 1970s by Eichten et al. [72] are a standard tool for the modeling of hadrons, firmly established in spectroscopic studies of heavy mesons in the charm and bottom sectors. The CQM is discussed in breadth elsewhere in this Encyclopedia [63], including also a concise derivation on QCD grounds.
@@
Figure 9: The light hadron spectrum of QCD as obtained by QCD. Horizontal lines and bands are the experimental values with their decay widths. QCD results are shown by solid circles. Vertical error bars represent combined statistical systematic error estimates. The pion ($\pi$), $S=-1$ kaon ($K$), and the $S=-2$ cascade baryon $\Xi$ have no error bars, because they were used to set the $u,d,s$ quark masses and the overall scale, respectively. The figure is adapted from Ref. [58].
@@
The CQM plays an important role in counting, ordering and classifying hadron states. For example, the CQM wave functions of baryons composed of three fermions are given by
$|\Phi_{3q}\rangle=|\psi_{color}\rangle\otimes|\psi_{flavor}\rangle|\psi_{spin}\rangle|\psi_{space}\rangle\>,$ (1)
@@
which in total must be antisymmetric. Color neutrality demands that the color-component is a singlet state which is antisymmetric by definition and construction. Thus, the remaining flavor-spin-space part must be symmetric which implies that the spin-flavor and the space part belong to the same symmetry class, i.e., both are symmetric or antisymmetric or of mixed symmetry. As explained in detail by Giannini [73] and in many textbooks, the symmetry condition imposes strict constraints on the allowed baryon configuration. In the $\{u,d,s\}$ sector, the spin and flavor degrees of freedom can be combined conveniently to form a SU(6) group. Combinatorics leads immediately to a total number of 216 allowed spin-flavor configurations. The respective irreducible representations are arranged into four SU(6) multiplets, given by a 20-plet of antisymmetric states, two 70-plets of mixed symmetry, and a 56-plet of symmetric states. The 20-plet and the 70-plets include flavor singlet states which typically are neglected because of lacking empirical evidence. The remaining irreducible components are flavor octets and decuplets, both of total spin $S=\frac{1}{2}$ and $S=\frac{3}{2}$, respectively. Evidently, a plethora of baryon states will be obtained by combining the SU(6) spin-flavor group structures and space parts provided by the O(3) group.
@@ 1 Counting Resonances
Soon it was realized that there was an obvious discrepancy between the number of resonances predicted by theory, as the rather successful constituent quark model, and those identified experimentally. Nathan Isgur (1947-2001) was probably the first who introduced the term missed resonances [74], thus honoring the mismatch between the more than 400 baryon states predicted group-theoretically by $SU(6)\times O(3)$ multiplet rulings (three 70-plets and four 56-plets) - about the same number is obtained by LQCD - and the number of experimentally verified resonances plus proton and neutron which amounts to a little more than 100 safely confirmed states PDG2024 [38]. The problem of missing resonances is a major issue of baryon spectroscopy. Final answers about the number of excited states of the nucleon and their spectral properties are still pending. Solutions are sought for both experimentally and theoretically.
@@
The current understanding of the internal quark-gluon composition of a hadron is illustrated in Fig. 8. The observed states, however, are the result of additional interactions with the set of meson-baryon scattering states of the same total quantum numbers. Hence, the complete picture of a hadron necessarily will lead to a multi-configuration problem, superimposing the internal, confined quark-gluon components and the external meson-baryon configurations, either as a virtual polarization cloud or as on-shell continuum channels. Obviously, quark models by themselves will account for such additional dynamical effects only in a weak coupling limit where mass shifts and other effects can be subsumed into effective phenomenologically determined parameters.
@@
In general, polarization effects, discussed below, induce self-energies which shift the bare QCD configurations in mass, and redistribute the spectroscopic strength i over the eigenstates of the coupled system. Formally, the eigenenergies of the interacting systems will be found to be moved into the complex plane where the imaginary parts define the decay and the lifetime of the polarized states. Thus, baryons heavier than the nucleon-pion system acquire rather short lifetimes because they may decay either by strong (and, to a minor degree, electromagnetic) interactions or, as in the case of hyperons, by weak interactions.
@@
The spin-isospin multiplet of $\Delta_{33}(1232)$ states, belonging to the $J^{P}=\frac{3}{2}^{+}$-baryon decuplet, is a typical case: As a QCD state, the Delta-resonance is understood as a spin-isospin vector ($\Delta S=1=\Delta I$) excitation of the nucleon by reorienting the spin and isospin of at least one quark, resulting in the $(J,I)=(\frac{3}{2}\frac{3}{2})$ configuration. That particular configuration is coupled, however, strongly to the continuum of pion-nucleon $P$-wave scattering states, which induces a rapid decay within a time span of $t_{1/2}\sim 10^{-23}~{}\mathrm{sec}$, corresponding to a spectral distribution of full width at half maximum (FWHM) $\Gamma_{\Delta}\simeq 120~{}\mathrm{MeV}$. In contrast to other baryon resonances, the Delta-resonance is prominently excited in practically all types of reactions, from virtual and real photo-excitation by electrons and photons to neutral and charged current neutrino-nucleon and neutrino-nucleus reactions, as well as in hadron scattering both on elementary proton targets and on nuclear targets. The description of those multi-channel and multi-configuration phenomena is the domain of the coupled channels methods discussed in the following sections.
@@ 2 Notations
The CQM spectroscopic notation is $|B{}^{2S+1}L_{J}\rangle_{\sigma}$ where $B=N,~{}\Delta$ denotes an octet or a decuplet, $L=S,~{}P,~{}D\ldots$ denotes the total orbital momentum, $|L-S|\leq J\leq L+S$ is the total angular momentum. The flavor symmetry character is specified by $\sigma=A,M,S$ for Asymmetric (A), Mixed (M), and Symmetric (S) .
@@
Traditionally, a different notation is used in meson-nucleon spectroscopy where states are classified by $L_{2I,2J}$, according to their pion-nucleon partial wave quantum number $L=S,~{}P,~{}D\ldots$, twice the total pion-nucleon isospin $I=\frac{1}{2},\frac{3}{2}$ and twice the total spin $J$. In that notation $I=\frac{1}{2}$, nucleon-like states are denoted by $P_{11}$, $P_{13}\ldots P_{12J}$, e.g., the nucleon is $P_{11}(940)$ and $P_{11}(1440)$, which is the Roper resonance (discovered in 1964 by L. David Roper (born 1935)). Thus, the resonance (pole) mass is added for distinguishing the resonances within the same partial wave.
@@
The non-strange $I=\frac{3}{2}$ decuplet states are labeled by $\Delta_{32J}$, where the best-known case is the lowest Delta-isobar resonance, $\Delta_{33}(1232)$ which was discovered by Enrico Fermi and his group in 1952. For pion-nucleon S-wave configurations and the higher partial waves $L\geq 2$, the standard spectroscopic notation $L_{2I,2J}$ is used, combined with the resonance mass. There are ambiguities in the choice of resonance masses. From a theoretical point of view, unstable, decaying states are characterized by the spectroscopic pole in the complex plane, including the full information on self-energies from the interactions with the surrounding continuum. However, frequently, the so-called Breit-Wigner mass (named after Gregory Breit (1899-1981) and Eugene Paul Wigner (1902-1995) [^11] Wigner received the Nobel Prize in Physics in 1963 for his contributions to the theory of the atomic nucleus and the elementary particles, particularly through the discovery and application of fundamental symmetry principles. ) is used instead [75].
@@
The international Particle Data Group (PDG), founded in 1957 by Murray Gell-Mann (1929-2019) and Arthur Hinton Rosenfeld (1926-2017) (the latter named “Godfather of Energy Efficiency”) [76]) uses a slightly different notation, as encountered in the latest PDG edition [38]: The proton and the neutron are distinguished by $p$, $I_{3}=+\frac{1}{2}$ and $n$, $I_{3}=-\frac{1}{2}$, thus accounting for the slight breaking of isospin symmetry in the nucleon iso- doublet as reflected in the p-n mass difference of about $1~{}\mathrm{MeV}$. Higher mass members of the $I=\frac{1}{2}$ nucleon family are denoted by $N(M_{n})J^{\pi}$, where $\pi=(-1)^{(}L+1)=\pm(1)$ is the parity of the state, composed of the orbital angular momentum $L$ of the partial wave and the intrinsic (negative) parity of the pion. $M_{n}$ is the resonance mass, also specifying the position of the state in the spectral sequence.
@@
The strangeness sector, $S=-1$ and $S=-3$ iso-singlet ($I=0$) baryons are denoted by $\Lambda(M_{k})J^{\pi}$ and $\Omega(M_{k})J^{\pi}$, respectively. Correspondingly, the iso-triplet ($I=1$) baryons of strangeness $S=-1$ and the iso-doublet ($I=\frac{1}{2}$) of strangeness $S=-2$ are classified by $\Sigma(M_{k})J^{\pi}$ and $\Xi(M_{k})J^{\pi}$, respectively. By historical reasons, the generic names Lambda ($\Lambda$), Sigma ($\Sigma^{0,\pm}$), Cascade ($\Xi^{0,-}$), and Omega ($\Omega^{-}$) are in use for the lowest octet and decuplet hyperons.
@@
By obvious reasons, most of the measurements are performed on proton (hydrogen) targets. Neutron data are obtained mainly on deuterium targets, used as a surrogate for non- existing elementary neutron targets and taking advantage of the loosely bound deuteron with a large distance between the constituents. The proton as a free particle was identified first by Ernest Rutherford (1871-1937) in 1920, but already in 1908 he was awarded the Nobel Prize in Chemistry for his investigations into the disintegration of the elements, and the chemistry of radioactive substances. The neutron was discovered by James Chadwick (1891-1974) in 1932 who received the Nobel Prize in Physics in 1935 for his discovery of the neutron.

### v2

@@ 2 From QCD to Hadrons
The, at first sight, seemingly simple picture of hadrons emerging from QCD is, in fact, deceptive. The complex field-theoretical dynamics lead to a large variety of spectroscopic configurations. The basic constituents, quarks and gluons, interact in a manifold manner, which will lead to a variety of intrinsic configurations forming the inner confined core of a hadron. A selection of expected elementary configurations is displayed in Fig. 8. The core region of a physical hadron will contain a mixture of those configurations, i.e., hadron wave functions are typically quantum mechanical superpositions of multi-quark, quark-gluon, gluon-gluon, etc. components - all embedded on top of a highly dynamical vacuum. The core region, albeit confined to an incredibly small volume of about $0.5\times 10^{-45}~\mathrm{m^{3}}$, is only part of the complete hadron. The quark-gluon core polarizes the vacuum and, as a consequence, is surrounded by a cloud of virtual particles, mainly consisting of pairs of pions and other mesons.
@@
Asymptotic freedom on one side and quark confinement on the other side are defining aspects of strong interaction physics. Gluon self-interactions and effects from the spontaneously broken chiral symmetry are important ingredients for constituent quark and hadron masses. The highly non-linear nature of QCD inhibits analytical or perturbative solutions for most parts of the accessible energy region, except for the highest energies, such as those reachable at the LHC at CERN. The mathematical complexities of QCD at energies relevant to hadron structure, spectroscopy, and interactions of hadrons can only be treated numerically. LQCD is the method of choice, successfully used to explore mesons and baryons through QCD principles.
@@ 1 QCD and Phenomenology
While the high-energy sector of QCD is well understood based on perturbative QCD, the low-energy limit remains a field with many open questions, reflecting the highly nontrivial task of handling a non-perturbative quantum field theory. Formally, the QCD path integral is the proper starting point. As discussed elsewhere in this encyclopedia, two alternative numerical schemes are currently pursued: In LQCD, a set of coupled correlation functions is derived from the path integral and propagated on a Euclidean (named after Euclid (about 325 — 265 BC), an ancient Greek mathematician) space-time grid, where the numerically produced energy levels [66]. One of the first successful LQCD descriptions of the mass spectrum of the octet mesons and the octet and decuplet baryons, respectively, was achieved by the Budapest-Wuppertal Collaboration [67, 68]. The mass spectrum is displayed in Fig. 9. Functional methods follow a different strategy by deriving sets of coupled Dyson-Schwinger equations (DSE) (named after Freeman John Dyson (1923-2020) and Julian Schwinger) by performing variational functional derivatives of the path integral [69].
@@
While LQCD and DSE approaches address hadron spectroscopy from the inner sector of the quark core of hadrons, effective field theories and phenomenological reaction models focus on the properties that reflect the nature of hadrons as dynamically generated composite states. Compositeness is another aspect of the dualism hidden in baryons and mesons, as observed in the resonances in scattering and production experiments. Until today, phenomenological approaches based on an underlying covariant field theory have been an important source of spectroscopic information on the quantum numbers of resonances, their formation, and their decay. The models are well-suited for large-scale numerical coupled-channel calculations using the partial-wave formalism. Meson production on the nucleon is described by the excitation of baryon resonances in photo- and meson-induced reactions. Baryon resonances are identified through the traces they leave as complex energy poles in the partial wave scattering matrix elements, which, in many cases, are strongly affected by coupled channel effects.
@@
At masses above the pion-nucleon threshold, baryon resonances must be regarded as superpositions of molecular-like, loosely bound, or unbound meson-nucleon configurations. They compete with the genuine QCD-type confined quark core configurations, as illustrated in Fig. 8. A given resonance will be composed of an arbitrary mixture of two building blocks. In some cases, the quark-type configurations will dominate; in others, the molecular configurations may prevail. An intensively and controversially debated candidate for the latter type of states is the $\Lambda^{\ast}(1405)$ (S=1S=-1) resonance, which falls outside the systematics of quark models but is characterized by a pronounced pole structure in the complex plane. The chiral unitary model [70] predicted a two-pole structure from a superposition of $\pi\Sigma$ and $\bar{K}N$ components. Years later, improved experimental and theoretical studies concluded that $\Lambda^{\ast}(1405)$ is dominantly a $\bar{K}N$ composite; however, they did not fully exclude $\pi\Sigma$ admixtures [71].
@@ 5 The Constituent Quark Model and Hadron Spectroscopy
Once QCD had established quarks and gluons as the constituents of hadrons, interest in nucleon spectroscopy was inevitable. However, in the early years of the QCD epoch, long before lattice QCD appeared on the horizon, the complexity of QCD dynamics did not allow for the explicit study of the non-perturbative regime of hadrons. However, in order to account for at least the most essential aspects, phenomenological models were developed. Bag models emphasize the confinement aspect; Skyrmion-inspired approaches concentrate on topological aspects, chirality, and soliton-like properties of hadrons. covariant energy density functionals, such as the still popular Nambu-Jona-Lasinio model, were derived, and the so-called sigma model, existing in linear and non-linear versions, has been extremely helpful in understanding spontaneous chiral symmetry breaking, the appearance of chiral condensates, and many other fundamental mechanisms inherent to hadrons. A detailed discussion, which goes beyond the purpose of this article, can be found in excellent topical monographs that address these topics in breadth.
Table 1: The mass spectrum of constituent quarks is used in the Constituent Quark Model. Masses are derived from the data of meson and baryon masses, such that the experimental masses are reproduced by the sum of constituent quark masses. The quark generations are indicated by separating lines. For further details, see Ref. [72].
@@
For spectroscopic purposes of hadrons, a model was required that could especially treat the quark degrees of freedom, because they were expected, for good reason, to prevail in the lowest baryon and meson resonances. Hence, oriented toward Gell-Mann’s former quark model, attempts started to add dynamics to the group-theoretical counting schemes. This was successful once gluons and other effects, such as polarization self-energies from the Dirac sea, were eliminated by ascribing their contributions to the effective masses of quarks—a standard method of quantum many-body theory, e.g., applied successfully in solid state and nuclear physics, and, as will be seen below, also in hadron spectroscopy. A working and solvable model was derived by replacing the bare QCD current quarks with effective constituent quarks. Constituent quarks incorporate the contributions to hadron spectra from the unresolved sectors of QCD. They can be understood as being surrounded by a polarization cloud of virtual excitations of the neglected degrees of freedom. A comparison of the current and constituent quark masses, Fig. 5 and Table 1, respectively, leads to the interesting conclusion that light quarks are affected the most by polarization effects, which may be interpreted as the polarization of the Dirac sea being important.
A still widely used CQM is the model of Isgur and Karl [73, 74, 75, 76, 77, 78, 79, 80]. It was an important step forward in understanding and predicting hadron spectra on phenomenological grounds. Mesons are described by quark-antiquark ($q\bar{q}$) configurations and baryons as states of three valence quarks (qqqqqq). The gluon component in hadron masses - actually accounting for a major part of the masses - is taken into account as sketched above, namely by assigning to the quarks effective constituent masses that are derived empirically from data. In Table 1, the CQM masses up to the top quark are listed.
@@
The constituent quarks interact through a Coulomb-like static potential [^11] Charles-Augustin de Coulomb (1736-1806) derived first the $1/r^{2}$ behavior of forces between electric charges, corresponding to a potential $\sim 1/r$., modeling gluon exchange, and are enclosed by a confinement potential, modeling color confinement. The confining potential, originally introduced in the early 1970s by Eichten et al. [81], known as the Cornell potential, is typically chosen as a power law in the radial coordinate. Spin- and flavor-dependent residual interactions are introduced in order to describe the fine structure of hadron spectra. Originally, the Cornell model was developed for heavy quarkonium spectroscopy. However, Godfrey and Isgur showed in the mid-1980s [82], that the CQM was indeed able to reproduce meson spectra surprisingly well, from the low-mass u,d,s multiplets to the region of heavy quarks. Modern versions of the Cornell potential and its off-springs are a standard tool for the modeling of hadrons, firmly established in spectroscopic studies of heavy mesons in the charm and bottom sectors. The CQM is discussed in breadth elsewhere in this Encyclopedia [72], including also a concise derivation on QCD grounds.
@@
Figure 9: The light hadron spectrum of QCD as obtained by QCD. Horizontal lines and bands are the experimental values with their decay widths. QCD results are shown by solid circles. Vertical error bars represent combined statistical systematic error estimates. The pion ($\pi$), S=1S=-1 kaon ($K$), and the S=2S=-2 cascade baryon $\Xi$ have no error bars, because they were used to set the $u,d,s$ quark masses and the overall scale, respectively. The figure is adapted from Ref. [83] using the LQCD results of Durr et al. [67].
@@
The CQM plays an important role in counting, ordering, and classifying hadron states. For example, the CQM wave functions of baryons composed of three fermions are given by
$|\Phi_{3q}\rangle=|\psi_{color}\rangle\otimes|\psi_{flavor}\rangle\otimes|\psi_{spin}\rangle\otimes|\psi_{space}\rangle\>,$ (1)
@@
which, in total, must be antisymmetric. Color neutrality demands that the color component be in a singlet state, which is antisymmetric by definition and construction. Thus, the remaining flavor-spin-space part must be symmetric, which implies that the spin-flavor and space parts belong to the same symmetry class, i.e., both are symmetric, antisymmetric, or of mixed symmetry. As explained in detail by Giannini [84] and in many textbooks, the symmetry condition imposes strict constraints on the allowed baryon configuration. In the $\{u,d,s\}$ sector, the spin and flavor degrees of freedom can be combined conveniently to form an SU(6) group. Combinatorics leads immediately to a total of 216 allowed spin-flavor configurations. The respective irreducible representations are arranged into four SU(6) multiplets, given by a 20-plet of antisymmetric states, two 70-plets of mixed symmetry, and a 56-plet of symmetric states. The 20-plet and the 70-plets include flavor singlet states, which are typically neglected due to a lack of empirical evidence. The remaining irreducible components are flavor octets and decuplets, both with total spin $S=\frac{1}{2}$ and $S=\frac{3}{2}$, respectively. Evidently, a plethora of baryon states will be obtained by combining the SU(6) spin-flavor group structures and space parts provided by the O(3) group.
@@ 1 Counting Resonances
Soon, it was realized that there was an obvious discrepancy between the number of resonances predicted by the rather successful constituent quark model and those identified experimentally. Nathan Isgur (1947-2001) was probably the first to introduce the term missing resonances [85], thus honoring the mismatch between the more than 400 baryon states predicted group-theoretically by $SU(6)\times O(3)$ multiplet rulings (three 70-plets and four 56-plets) - about the same number is obtained by LQCD - and the number of experimentally verified resonances plus proton and neutron, which amounts to a little more than 100 safely confirmed states PDG2024 [41]. The problem of missing resonances is a major issue in baryon spectroscopy. Final answers regarding the number of excited states of the nucleon and their spectral properties are still pending. Solutions are sought both experimentally and theoretically.
@@
The current understanding of the internal quark-gluon composition of a hadron is illustrated in Fig. 8. The observed states, however, are the result of additional interactions with the set of meson-baryon scattering states that possess the same total quantum numbers. Hence, the complete picture of a hadron will necessarily lead to a multi-configuration problem, superimposing the internal, confined quark-gluon components and the external meson-baryon configurations, either as a virtual polarization cloud or as on-shell continuum channels. Obviously, quark models by themselves will account for such additional dynamical effects only in a weak coupling limit, where mass shifts and other effects can be subsumed into effective, phenomenologically determined parameters.
@@
In general, polarization effects, discussed below, induce self-energies that shift the bare QCD configurations in mass and redistribute the spectroscopic strength over the eigenstates of the coupled system. Formally, the eigenenergies of the interacting systems will be found to have moved into the complex plane, where the imaginary parts define the decay and the lifetime of the polarized states. Thus, baryons heavier than the nucleon-pion system acquire rather short lifetimes because they may decay either by strong (and, to a minor degree, electromagnetic) interactions or, as in the case of hyperons, by weak interactions.
@@
The spin-isospin multiplet of $\Delta_{33}(1232)$ states, belonging to the $J^{P}=\frac{3}{2}^{+}$-baryon decuplet, is a typical case: As a QCD state, the Delta-resonance is understood as a spin-isospin vector ($\Delta S=1=\Delta I$) excitation of the nucleon by reorienting the spin and isospin of at least one quark, resulting in the $(J,I)=(\frac{3}{2}\frac{3}{2})$ configuration. That particular configuration is coupled, however, strongly to the continuum of pion-nucleon $P$-wave scattering states, which induces a rapid decay within a time span of $t_{1/2}\sim 10^{-23}~\mathrm{sec}$, corresponding to a spectral distribution of full width at half maximum (FWHM) $\Gamma_{\Delta}\simeq 120~\mathrm{MeV}$. In contrast to other baryon resonances, the Delta-resonance is prominently excited in practically all types of reactions, from virtual and real photo-excitation by electrons and photons to neutral and charged current neutrino-nucleon and neutrino-nucleus reactions, as well as in hadron scattering on both elementary proton targets and nuclear targets. The description of those multi-channel and multi-configuration phenomena is the domain of the coupled channels methods discussed in the following sections.
@@ 2 Notations
The CQM spectroscopic notation is $|B{}^{2S+1}L_{J}\rangle_{\sigma}$, where $B=N,~\Delta$ denotes an octet or a decuplet, $L=S,~P,~D\ldots$ denotes the total orbital momentum, and $|L-S|\leq J\leq L+S$ is the total angular momentum. The flavor symmetry character is specified by $\sigma=A,M,S$ for Asymmetric (A), Mixed (M), and Symmetric (S).
@@
Traditionally, a different notation is used in meson-nucleon spectroscopy, where states are classified by $L_{2I,2J}$ according to their pion-nucleon partial wave quantum number $L=S,~P,~D\ldots$, twice the total pion-nucleon isospin $I=\frac{1}{2},\frac{3}{2}$, and twice the total spin $J$. In that notation $I=\frac{1}{2}$, nucleon-like states are denoted by $P_{11}$, $P_{13}\ldots P_{12J}$, e.g., the nucleon is $P_{11}(940)$ and $P_{11}(1440)$, which is the Roper resonance (discovered in 1964 by L. David Roper (born 1935)). Thus, the resonance (pole) mass is added to distinguish the resonances within the same partial wave.
@@
The non-strange $I=\frac{3}{2}$ decuplet states are labeled by $\Delta_{3/2,J}$, where the best-known case is the lowest Delta-isobar resonance, $\Delta_{33}(1232)$, which was discovered by Enrico Fermi and his group in 1952. For pion-nucleon S-wave configurations and the higher partial waves $L\geq 2$, the standard spectroscopic notation $L_{2I,2J}$ is used, combined with the resonance mass. There are ambiguities in the choice of resonance masses. From a theoretical point of view, unstable, decaying states are characterized by the spectroscopic pole in the complex plane, which includes full information on self-energies from the interactions with the surrounding continuum. However, frequently, the so-called Breit-Wigner mass (named after Gregory Breit (1899-1981) and Eugene Paul Wigner (1902-1995) [^12] Wigner received the Nobel Prize in Physics in 1963 for his contributions to the theory of the atomic nucleus and elementary particles, particularly through the discovery and application of fundamental symmetry principles. ) is used instead [86].
@@
The international Particle Data Group (PDG), founded in 1957 by Murray Gell-Mann (1929-2019) and Arthur Hinton Rosenfeld (1926-2017) (the latter named “Godfather of Energy Efficiency”) [87]) uses a slightly different notation, as encountered in the latest PDG edition [41]: The proton and the neutron are distinguished by $p$, $I_{3}=+\frac{1}{2}$ and $n$, $I_{3}=-\frac{1}{2}$, thus accounting for the slight breaking of isospin symmetry in the nucleon iso-doublet, as reflected in the p-n mass difference of about $1~\mathrm{MeV}$. Higher mass members of the $I=\frac{1}{2}$ nucleon family are denoted by $N(M_{n})J^{\pi}$, where $\pi=(-1)^{L+1}=\pm 1$ is the parity of the state, composed of the orbital angular momentum $L$ of the partial wave and the intrinsic (negative) parity of the pion. $M_{n}$ is the resonance mass, also specifying the position of the state in the spectral sequence.
@@
The strangeness sector, S=1S=-1 and S=3S=-3 iso-singlet ($I=0$) baryons are denoted by $\Lambda(M_{k})J^{\pi}$ and $\Omega(M_{k})J^{\pi}$, respectively. Correspondingly, the iso-triplet ($I=1$) baryons of strangeness S=1S=-1 and the iso-doublet ($I=\frac{1}{2}$) of strangeness S=2S=-2 are classified by $\Sigma(M_{k})J^{\pi}$ and $\Xi(M_{k})J^{\pi}$, respectively. By historical reasons, the generic names Lambda ($\Lambda$), Sigma ($\Sigma^{0,\pm}$), Cascade ($\Xi^{0,-}$), and Omega ($\Omega^{-}$) are used for the lowest octet and decuplet hyperons.
@@
For obvious reasons, most of the measurements are performed on proton (hydrogen) targets. Neutron data are obtained mainly from deuterium targets, which are used as a surrogate for non-existent elementary neutron targets, taking advantage of the loosely bound deuteron with a large distance between its constituents. The proton, as a free particle, was identified first by Ernest Rutherford (1871-1937) in 1920. However, already in 1908, he was awarded the Nobel Prize in Chemistry for his investigations into the disintegration of the elements and the chemistry of radioactive substances. The neutron was discovered by James Chadwick (1891-1974) in 1932, who received the Nobel Prize in Physics in 1935 for his discovery of the neutron.